Fibre Damage Friday 27th December 2019 15:24:00


We have been alerted to what appears to be a repeat of the malicious attack suffered on 12th September. Initial reports are looking like fire damage to the same fibre junction as before but we are still investigating the extent of the damage.

We aim to have this resolved as quickly as possible and will provide updates as we have them.

Apologies to all affected customers.

All but one customer's connectivity is back online.

We are doing all we can with the final affected line and are in direct contact with said customer.

We apologise for any issues this has caused; some people evidently have more flammable spirit than festive spirit. We hope this incident has not dampened your festive celebrations.

All splicing works have been completed.

We are undergoing some final testing and expect all to be as planned, however if you are still experiencing any issues please contact us.

Restorative work is still continuing, however the bulk of the damage has been repaired.

Engineers are still on site completing the work and major changes will continue to be notified here.

If you are having connection issues please do advise us ASAP.

Splicing work continued through the night with two of three 96F cables fully replaced and the third nearing completion.

Many connections came back online by 4am.

We are currently investigating continuity on the original fault fibres and will update again soon.

The 96F cable has been replaced and splicing continues onto the next cable allocation.

Work will continue through the night but the next update will be later this morning.

Splicing has now begun on the two 48f sections of fibre, all 96f will be spliced before continuing to the next section of fibre.

Further updates will continue as work is progressed

The first fibre has now been pulled through and the preparation of the joints is underway prior to the splicing work being undertaken.

Work is to continue throughout the night.

Fibre is continued to be spliced, however, due to the large amount of damage all the fibre has to be replaced. Engineers will continue to work throughout the night to progress the repair.

Engineers are confirmed as being on site and are making good progress, no estimated time of resolution has been confirmed yet however they will continue to keep us informed.

Current update: - Services hard down due to vandalism

Action: - Engineers attending from various locations, arriving between 1730-1900

Next Step: - Await updates from engineering team and provide updates as necessary